OpenStudy (anonymous):

find the volume of a hemisphere with a diameter of 17 cm showing all work .... help ! :/

OpenStudy (anonymous):

radius is 8.5 then Find the volume of a whole sphere with the given radius and divide that by 2

OpenStudy (anonymous):

Hemisphere is half of a sphere....do you know the formula for the volume of a sphere

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

Nice pic....but we want half of it

OpenStudy (unklerhaukus):

\[V_{\text{Sphere}}=\frac{4}{3}\pi r^3\] \[V_{\text{Hemiphere}}=\frac12\times\frac{4}{3}\pi r^3\] \[D=2r\]

OpenStudy (unklerhaukus):

\[r^3=\left(\frac{D}{2}\right)^3=\frac{D^3}{2^3}=\frac{D^3}{8}\]

OpenStudy (unklerhaukus):

\[V_{\text{Hemiphere}}=\frac12\times\frac{4}{3}\times\pi \times\frac{17^3}8\]

OpenStudy (unklerhaukus):